### v1

# DHG-Bench: A Comprehensive Benchmark on Deep Hypergraph Learning

Abstract
Although conventional deep graph models have achieved great success in relational learning, their focus on pairwise relationships limits their capacity to learn pervasive higher-order interactions in real-world complex systems, which can be naturally modeled as hypergraphs. To tackle this, hypergraph neural networks (HNNs), the dominant approach in deep hypergraph learning (DHGL), has garnered substantial attention in recent years. Despite the proposal of numerous HNN methods, there is no comprehensive benchmark for HNNs, which creates a great obstacle to understanding the progress of DHGL in several aspects: (i) insufficient coverage of datasets, algorithms, and tasks; (ii) a narrow evaluation of algorithm performance; and (iii) inconsistent dataset usage, preprocessing, and experimental setups that hinder comparability. To fill the gap, we introduce DHG-Bench, the first comprehensive benchmark for DHGL. Specifically, DHG-Bench integrates 20 diverse datasets spanning node-, edge-, and graph-level tasks, along with 16 state-of-the-art HNN algorithms, under consistent data processing and experimental protocols. Our benchmark systematically investigates the characteristics of HNNs in terms of four dimensions: effectiveness, efficiency, robustness, and fairness. Further, to facilitate reproducible research, we have developed an easy-to-use library for training and evaluating different HNN methods. Extensive experiments conducted with DHG-Bench reveal both the strengths and inherent limitations of existing algorithms, offering valuable insights and directions for future research. The code is publicly available at: https://github.com/Coco-Hut/DHG-Bench.
1 Introduction
Graphs are powerful tools for capturing structural and relational properties in various domains, such as social networks (Fan et al., 2019), e-commerce networks (Liu et al., 2021), and financial networks (Li et al., 2024b). Graph Neural Networks (GNNs) have emerged as the dominant approach for learning such data, owing to their strong ability to capture informative topological representations. However, many real-world systems involve multi-way or group-wise interactions beyond the pairwise connections of graphs. For instance, multiple authors co-write a paper in co-authorship networks (Yang et al., 2022), and groups of proteins interact collectively in biological systems (Kim et al., 2024b). These higher-order interactions are naturally modeled by hypergraphs, where each hyperedge connects an arbitrary number of nodes. Directly modeling such higher-order relations has shown improvements across various machine learning tasks compared to graph-based models (Benson et al., 2016; Li & Milenkovic, 2017; Srinivasan et al., 2021; Wang et al., 2023b). This motivates the generalization of GNNs to hypergraphs.
Building on this need, deep hypergraph learning (DHGL) has been developed for learning higher-order knowledge, with hypergraph neural networks (HNNs) emerging as its principal methodology and attracting considerable research interest in recent years. These methods target using neural functions to transform given nodes, hyperedges, and their features into vector representations (i.e., embeddings) that effectively preserve higher-order information. HNNs have shown state-of-the-art performance in industrial and scientific applications, including product recommendation (Ji et al., 2020), 3D object detection (Gao et al., 2012), disease prediction (Gao et al., 2015), and more (Tian et al., 2009; Xu et al., 2018; Zhao et al., 2023).
Despite the emerging studies of HNN algorithms, there lacks a comprehensive and unified benchamrk. HyFER (Hwang et al., 2021) is a modular, easy-to-use framework that supports efficient implementation and fair evaluation of HNNs across diverse datasets, models, and tasks. THU-DH (Gao et al., 2022) is an open-source PyTorch-based toolbox designed for general HNNs. It supports various hypergraph pre-processing methods (e.g., sampling, expansion) and convolution operators, facilitating the evaluation of HNNs. However, existing DHGL benchmarks still suffer from several limitations: (i) The scope of the datasets, algorithms, and benchmark tasks considered in existing studies remains insufficiently comprehensive. For example, HyFER integrates only three HNN methods, while THU-DH supports only approaches published before 2022, thereby excluding many recent state-of-the-art methods. Moreover, the employed datasets are predominantly node-level homophilic hypergraphs, overlooking both heterophilic hypergraphs and those designed for hypergraph-level tasks. This narrow coverage restricts the evaluation of model generalizability across diverse structural patterns. (ii) Algorithm evaluation in existing studies is largely limited to the effectiveness dimension, offering an incomplete picture of model performance and characteristics. Beyond effectiveness, investigating the efficiency and complexity of the algorithm is imperative, yet often overlooked in the literature. Moreover, the trustworthiness of algorithms, particularly their robustness and fairness, remains under-explored despite being crucial for determining their suitability in industrial applications. (iii) The use of different benchmark datasets, data processing approaches, and experimental setups (e.g., parameter settings, initialization procedures) hinders reproducibility and fair comparison across studies. Hence, there is an urgent necessity within the community to develop a standardized and comprehensive benchmark for DHGL.
To bridge this gap, we introduce DHG-Bench, which serves the first open-sourced and comprehensive benchmark for HNNs. First, DHG-Bench implements a wide range of HNN algorithms through unified APIs, while also adopting consistent data processing and data splitting approahces for fair comparisons. Our benchmark encompasses 16 state-of-the-art HNN algorithms and 20 diverse hypergraph datasets covering node-level, edge-level, and graph-level tasks. Moreover, we systematically investigate the characteristics of HNNs in terms of four dimensions: effectiveness, efficiency, robustness, and fairness. Based on these four dimensions, we conduct an extensive comparative study of existing HNN algorithms in different scenarios. For effectiveness, DHG-Bench offers a fair and comprehensive comparison of HNN algorithms on node- and edge-level tasks involving both homophilic and heterophilic hypergraphs. It is also the first benchmark to incorporate the hypergraph-level classification task into evaluation, enabling a more complete assessment of model capabilities across different task levels. For efficiency, DHG-Bench empirically evaluates the efficiency of representative HNNs in terms of both time and space complexity. For robustness, DHG-Bench evaluates HNN models under perturbations at three levels: hypergraph structure, node features, and supervision signals. For fairness, DHG-Bench evaluates the demographic fairness of existing HNNs through a fair node classification task, revealing their performance disparities across different sensitive groups.
Through extensive experiments, we have the following key insights: (1) HNN models show substantial performance variability across different datasets and tasks, and recent approaches have yet to demonstrate consistently better generalization across them. (2) While most HNN methods struggle to strike a satisfactory balance among predictive performance, efficiency, and scalability, TFHNN, by decoupling message passing from model training, achieves the most favorable trade-off. (3) HNN algorithms are susceptible to data perturbations from different aspects, particularly from supervision signals. (4) Although HNNs achieve improved predictive performance, they suffer from significant fairness issues. Our main contributions are summarized as follows:
DHG-Bench enables a fair and unified comparison among 16 state-of-the-art HNN algorithms and 20 diverse hypergraph datasets covering node-level, edge-level, and graph-level tasks. To the best of our knowledge, this is the first comprehensive benchmark for deep hypergraph learning.
We conduct a systematic analysis of HNN methods from various dimensions, including effectiveness, efficiency, robustness, and fairness. Through extensive experiments, we reveal both the strengths and limitations of existing algorithms, offering valuable insights to inform and inspire future research in this field.
We release DHG-Bench, an easy-to-use open-source benchmark library to support future DHGL research. Besides, with our toolkit, users can readily evaluate their algorithms or datasets with less effort.
2 Preliminary
In this section, we first introduce the notations used throughout the paper, and then give the concepts of three benchmark tasks spanning node-, hyperedge-, and hypergraph-level prediction.
Notations. Let represent a hypergraph with vertex set = and hyperedge set = . denotes the dimensional node feature matrix. As each hyperedge is a subset of , we denote as the size of .
Node Classification. Given the labelded node set and their labels , where each node is associated with a label belonging to one of classes. Semi-supervised node classification aims to train a node classifier to predict labels of the remaining nodes .
Hyperedge Prediction. Given a hypergraph , we deonte as the target set which typically consists of (a) unobserved hyperedges or (b) new hyperedges that will arrive in the near future. Each element in is referred to as a hyperedge candidate, denoted by , as it may belong to . Hyperedge prediction aims to train a hyperedge classifier to preidct whether candidate belongs to the target set or not.
Hypergraph Classification. Let as the hypergraph set. Given the labeled hypergraph set and their labels , where each hypergraph is assigned a label . Semi-supervised hypergaph classification task aims to train a hypergraph classifier to predict labels of the unlabeled hypergraphs .
3 Benchmark Design
In this section, we introduce the DHG-Bench in terms of datasets (Section 3.1), algorithms (Section 3.2), and research questions (Section 3.3) that guide the benchmark design. The overview of DHG-Bench is illustrated in Figure 1.
3.1 Benchmark Datasets
To comprehensively and effectively evaluate the characteristics of various HNN algorithms, we have integrated 20 real-world datasets from various domains for node-level, hyperedge-level, and hypergraph-level tasks. We briefly introduce each category in this section and provide detailed descriptions in Appendix A.1.
Node-level Classification Datasets. For the most prominent task in hypergraph learning, node classification, we utilize 11 hypergraph datasets spanning diverse domains and exhibiting varying characteristics. As summarized in Table A1 in Appendix A, the benchmark includes 6 homophilic datasets: 2 co-citation networks (Cora, Pubmed (Yadati et al., 2019)), 2 co-authorship networks (Cora-CA, DBLP (Yadati et al., 2019)), and 2 large-scale benchmarks (Walmart (Chien et al., 2022), and Trivago (Kim et al., 2023)) that model co-purchasing, and co-click interactions, respectively. In addition, we incorporate 5 heterophilic datasets that capture multi-way relationships among entities in various contexts, including two information networks (Actor (Li et al., 2025) and Yelp (Chien et al., 2022)), an e-commerce network (Amazon-ratings (Li et al., 2025)), and two social networks (Twitch-gamers, Pokec (Li et al., 2025)). To further investigate potential fairness concerns in existing HNN algorithms, we include three fairness-sensitive datasets (German, Bail, and Credit (Wei et al., 2022)), which contain sensitive node attributes such as gender, race, and age. Detailed statistics of these datasets are presented in Table A2. In line with previous studies (Feng et al., 2019; Chien et al., 2022; Wang et al., 2023a; Tang et al., 2025), we adopt a 50%/25%/25% split for training, validation, and testing in the node classification task.
Hyperedge-level Prediction Datasets. For the hypergraph prediction task, we consider 4 widely used real-world datasets: two co-citation networks (Cora, Pubmed) and two co-authorship networks (Cora-CA, DBLP-CA), as adopted in prior studies (Hwang et al., 2022; Ko et al., 2025). To support a more comprehensive evaluation, we further include two heterophilic datasets, Actor and Pokec, characterized by low hyperedge homophily (Li et al., 2025). Following the protocol in (Hwang et al., 2022; Ko et al., 2025; Yu et al., 2025), we randomly split the hyperedges (i.e., positive samples) into training (60%), validation (20%), and test (20%) sets, repeated over five random seeds. To rigorously assess the performance of various HNNs, we construct four different validation and test sets, each of which has different negative examples with various degrees of difficulty. Specifically, for each validation and test set, we sample an equal number of negative examples as the positive ones. Following [1], we adopt a mixed negative sampling (NS) strategy that incorporates three widely adopted heuristic NS methods: sized NS (SNS), motif NS (MNS), and clique NS (CNS) (Patil et al., 2020), thereby increasing the diversity of negative samples.
Hypergraph-level Classification Datasets. For the hypergraph classification task, We consider 6 benchmark datasets introduced in a recent study (Feng et al., 2024), as summarized in Table A3. RHG-10 is a synthetic datasets which encompass ten distinct hypergraph structure factors (e.g., Hyper Flower and Hyper Pyramid). RHG-3 is a synthetic dataset consisting of randomly generated hypergraphs with three distinctively various hypergraph structures: hyper-pyramid, hyper-check-table, and hyper-wheel. IMDB-Dir-Form and IMDB-Dir-Genre are two datasets constructed by the co-director relationship from the original IMDB dataset 111https://www.imdb.com/. Steam-Player is a player-based dataset, where each hypergraph captures tag co-occurrence relationships among games played by a user. Twitter-Friend is a social media dataset, where each hypergraph represents the friendship network of a specific Twitter user. For hypergraph classification, we adopt a 80%/10%/10% train/validation/test data split.
3.2 Benchmark Algorithms
We have integrated 16 representative and competitive HNN algorithms in DHG-Bench. They are divided into three technique categories: Spectral-based HNN, Spatial-based HNN, and Tensor-based HNN. We briefly introduce each category in the following, and more details are provided in Appendix A.2.
Spectral-based HNN Algorithms. Spectral-based HNN methods are a class of algorithms that leverage the spectral theory of hypergraph Laplacians to perform message propagation and feature transformation. These models operate in the spectral domain, where convolution operations are defined based on the eigendecomposition of a hypergraph Laplacian matrix derived from the adjacency matrix of the hypergraph (Saxena et al., 2024; Wang et al., 2024a). We integrate 10 representative spectral-based algorithms including HGNN (Feng et al., 2019), HyperGCN (Yadati et al., 2019), HCHA (Bai et al., 2021), LEGCN (Yang et al., 2022), HyperND (Prokopchik et al., 2022), PhenomNN (Wang et al., 2023b), SheafHyperGNN (Duta et al., 2023), HJRL (Yan et al., 2024), DPHGNN (Saxena et al., 2024), and TF-HNN (Tang et al., 2025).
Spatial-based HNN Algorithms. In contrast to spectral-based method, spatial-based methods focus on the local connectivity of each node without going to the spectral domain (Wang et al., 2024a). They commonly apply a two-stage neighborhood aggregation scheme to learn hypergraph representations, which entails updating the hyperedge embedding by aggregating embeddings of its incident nodes, and updating the node representation by propagating information from representations of its incident hyperedges (Gao et al., 2022; Wang et al., 2024a). We consider 4 spatial-based algorithms including HNHN (Dong et al., 2020), UniGNN (Huang & Yang, 2021), AllSetTransformer (Chien et al., 2022), and ED-HNN (Wang et al., 2023a). Since multiple UniGNN variants exist (e.g., UniGAT, UniGIN, UniGCNII), we report only the results of the most competitive variant, UniGCNII, for simplicity. We also note that our open-source library supports the implementations of other variants.
Tensor-based HNN Algorithms. These methods leverage tensor operations that provide a structured and effective means of capturing the complexity of hypergraph interactions (Wang et al., 2024b; 2025). In our DHG-Bench, we integrate two representative tensor-based algorithms: EHNN (Kim et al., 2022), and T-HyperGNN (Wang et al., 2024a).
3.3 Research Questions
We systematically design the DHG-Bench to comprehensively evaluate the existing HNN algorithms and inspire future research. In particular, we aim to investigate the following research questiones.
RQ1: How effective are the existing HNN algorithms on different benchmark tasks?
Motivation and Experiment Design. Previous research on HNNs has been hindered by inconsistent data preprocessing, varying hyperparameter settings, and limited benchmark tasks, making it difficult to fairly and comprehensively evaluate the performance of different methods. Given the standardized experimental environment provided by DHG-Bench, our first research question is to revisit the effectiveness of existing HNN algorithms. By addressing this question, we aim to gain a deeper understanding of their strengths and limitations, and to identify directions for potential improvement. To this end, we evaluate the performance of HNN methods on 17 widely used hypergraph datasets across three representative tasks: node classification, hyperedge prediction, and hypergraph classification. Detailed experimental settings are provided in Appendix B, and evaluation results are presented in Section 4.1.
RQ2: How effiecient are these HNN algorithms in terms of time and space?
Motivation and Experiment Design. As existing HNN algorithms often rely on the construction of various hypergraph Laplacians or adopt a two-stage message passing scheme during training, they can naturally introduce additional computational and memory overhead compared to standard message passing GNNs. However, the efficiency and scalability of HNN methods have been largely overlooked in prior work, despite being critical for understanding the trade-off between computational cost and task performance. A thorough evaluation of these aspects is essential for assessing the practical suitability of each method for real-time and large-scale deployment. To answer this question, we evaluate the efficiency of each HNN method in terms of both time and space. Specifically, we record the total training time required to reach the best validation performance, as well as the peak GPU memory usage during training. The evaluation resutls are repo in Sec 4.2.
RQ3: How robust are the HNN algorithms under various forms of data perturbation?
Motivation and Experiment Design. Robustness evaluation is crucial for assessing the reliability of HNN algorithms under real-world conditions, especially in high-stakes domains such as healthcare and finance. Ideally, a robust HNN should maintain stable performance when exposed to various data perturbations (e.g., random noise and adversarial attacks). Such evaluation not only reveals potential vulnerabilities in existing methods but also guides the development of more resilient HNN models. To answer this research question, we conduct a systematic robustness evaluation from three perspectives: structure-level robustness, feature-level robustness, and supervision-level robustness. For structure robustness, we randomly perturb the hypergraph by adding or removing a proportion of node–hyperedge incidence connections. For feature robustness, we apply two types of random perturbations: (1) masking a certain proportion of node features, and (2) injecting Gaussian noise based on the dimensionality of node features. Furthermore, we assess the resilience of HNN algorithms to two types of supervision-level perturbations: (1) varying the sparsity of the supervision signal by randomly masking a proportion of training labels, and (2) introducing label noise by randomly assigning incorrect labels to a subset of the training samples. The evaluation results are shown in Sec 4.3.
RQ4: To what extent do existing HNN algorithms induce unfairness across demographic groups?
Motivation and Experiment Design. Fairness has recently emerged as a critical concern in graph machine learning (GML) (Dong et al., 2023). Prior studies have shown that representations learned by GNNs can result in biased predictions, often favoring certain demographic groups defined by sensitive attributes (e.g., gender, race) (Ling et al., 2023; Zhu et al., 2024; Yang et al., 2024). Such bias hinders the deployment of GML models in high-stakes applications such as crime prediction (Suresh & Guttag, 2019) and credit evaluation (Yeh & Lien, 2009). Despite its importance, fairness in deep hypergraph learning has received little attention. To the best of our knowledge, this work presents the first benchmark evaluation of fairness in this context, which is crucial for developing ethically sound and reliable HNN models. To answer this question, we conduct node classification on three fairness-sensitive datasets (German, Bail, and Credit (Wei et al., 2022)), each of which contains demographic-sensitive attributes. We assess the fairness of various HNN methods using two widely adopted group fairness metrics: demographic parity () (Dwork et al., 2012), and equalized odds () (Hardt et al., 2016). Please refer to Appendix A.1 and Appendix B.2 for detailed informatoin about the datasets and metrics, respectively. The fairness evaluation results are reported in Sec 4.4.
4 Experiment Results and Analysis
4.1 Effectiveness Evaluation (RQ1)
| Method | Cora | Pubmed | Cora-CA | DBLP-CA | Yelp | Warmart | Trivago | Actor | Ratings | Gamers | Pokec |
| HGNN | 77.901.17 | 86.170.52 | 82.840.46 | 91.000.27 | 33.710.24 | 77.120.12 | 57.671.61 | 77.830.37 | 28.050.28 | 52.380.56 | 57.870.76 |
| HyperGCN | 78.381.63 | 87.420.42 | 81.651.58 | 89.510.18 | 29.290.55 | 68.750.56 | 42.391.25 | 81.820.39 | 27.340.72 | 51.320.72 | 57.510.54 |
| HCHA | 77.841.23 | 86.330.54 | 83.010.58 | 91.180.30 | 33.130.23 | 77.660.18 | 52.503.43 | 78.300.47 | 28.330.34 | 52.350.71 | 58.190.45 |
| LEGCN | 74.361.03 | 87.520.50 | 74.591.04 | 85.160.14 | OOM | 62.980.09 | 33.451.45 | 85.340.45 | 28.210.50 | 51.310.65 | 59.660.63 |
| HyperND | 79.230.63 | 86.730.56 | 83.190.71 | 91.340.19 | OOM | 75.100.54 | 87.191.89 | 83.190.92 | 28.320.38 | 52.390.60 | 57.651.08 |
| PhenomNN | 78.971.41 | 87.810.12 | 84.051.05 | 91.830.25 | OOM | OOM | OOM | 83.140.49 | 28.490.41 | 51.800.73 | 58.430.92 |
| SheafHyperGNN | 79.030.90 | 87.100.47 | 84.080.50 | 91.090.31 | OOM | OOM | OOM | 85.000.32 | 28.350.57 | 52.070.53 | 59.060.37 |
| HJRL | 78.671.47 | 87.980.49 | 83.720.74 | OOM | OOM | OOM | OOM | 71.540.64 | 26.900.55 | 51.620.61 | 57.570.47 |
| DPHGNN | 76.401.36 | 86.720.33 | 82.131.13 | OOM | OOM | OOM | OOM | 83.650.59 | 28.571.07 | 52.360.59 | 58.200.58 |
| TF-HNN | 79.471.31 | 87.900.37 | 84.190.89 | 91.380.24 | 35.160.54 | 77.040.12 | 90.790.79 | 85.960.41 | 28.560.68 | 52.340.53 | 59.170.52 |
| HNHN | 75.241.38 | 85.661.28 | 76.511.34 | 85.840.07 | 25.860.63 | 65.210.28 | 53.751.43 | 81.200.36 | 27.290.70 | 51.120.65 | 58.550.93 |
| UniGNN | 79.411.24 | 87.570.54 | 83.491.58 | 91.710.20 | 31.090.61 | 76.260.58 | 36.150.56 | 84.610.46 | 28.390.64 | 52.500.57 | 58.560.73 |
| AllSetTransformer | 78.021.43 | 87.790.30 | 82.950.62 | 91.510.22 | 33.180.88 | 78.610.13 | 59.924.02 | 85.660.41 | 27.321.11 | 51.740.75 | 58.550.56 |
| ED-HNN | 78.580.52 | 87.650.23 | 82.980.93 | 91.550.19 | 34.840.93 | 77.900.21 | 75.992.60 | 85.770.46 | 28.380.31 | 50.540.23 | 58.680.40 |
| EHNN | 76.511.52 | 87.120.31 | 81.680.81 | 90.470.43 | 34.093.19 | 77.950.14 | OOM | 86.210.49 | 28.950.81 | 52.140.76 | 58.231.07 |
| T-HyperGNN | 74.201.37 | 86.280.62 | 75.011.44 | 85.440.14 | OOM | 73.480.33 | OOM | 85.320.48 | 24.631.22 | 51.820.38 | 58.820.49 |
To investigate the progress and effective ranges of existing HNN methods, we compare their performance across benchmark tasks at the node, edge, and graph levels.
4.1.1 Effectiveness on node classification task. We conduct node classification experiments on 11 datasets spanning both homophilic and heterophilic hypergraphs of varying scales. As shown in Table 1, we have the following key observations: (1) TFHNN consistently ranks among the top-performing methods across diverse datasets Specifically, it achieves the best or second-best results on 8 out of 11 datasets and delivers prediction accuracy comparable to the top-performing models on the remaining three. These results demonstrate the effectiveness and robustness of its training-free message-passing module and decoupled architecture in the node classification task. (2) Existing HNN methods generally achieve lower node classification performance on heterophilic datasets than on homophilic ones. For instance, on four homophilic academic networks, all methods consistently achieve accuracies above 74%, whereas on the heterophilic Yelp and Ratings datasets, the highest accuracy is below 36%. The results indicate that heterophilic patterns can impair the quality of hypergraph representation learning. (3) Most existing HNN algorithms either face scalability issues or experience significant accuracy degradation on large-scale datasets. On the Trivago dataset, for example, six methods encounter OOM errors, and eight of the remaining ten achieve accuracies below 80%, with at least a 14.8% gap from TFHNN, the best-performing model on this dataset.
4.1.2 Effectiveness on hyperedge prediction task. For the edge-level task, we conduct experiments on six widely used real-world datasets, reporting AUROC and AP (Average Precision). As shown in Table 2, we find that: (1) Recent proposed HNN methods that achieve superior node classification performance fail to remain equally competitive in hyperedge prediction. In particular, we find that HGNN, HyperGCN, and EHNN, all proposed before 2022, together account for the best and second-best performance on the six datasets for the edge-level task. In contrast, more recent HNNs (e.g., HJRL, TFHNN, and DPHGNN) often show a notable performance gap compared to these three. For example, on Cora-CA, DPHGNN achieves an AUROC of 82.89% and an AP of 83.78%, representing decreases of 9.84% and 9.64%, respectively, relative to HyperGCN. (2) The performance gaps between most HNN algorithms and the corresponding state-of-the-art models vary greatly across datasets, with the magnitude of the gap heavily dependent on the dataset. For instance, on the Pubmed dataset, PhenomNN, UniGNN, and AllSetTransformer each achieve an AUROC within 5.5% of the state-of-the-art EHNN (79.54%). In contrast, on the Actor dataset, their AUROC scores fall at least 15% below that of the best-performing model, HGNN. This suggests that current research lacks HNN methods with strong generalization ability for hyperedge prediction across diverse datasets.
4.1.3 Effectiveness on hypergraph classification task. We conduct hypergraph classification task on two synthetic datasets and four real-world datasets. As current tensor-based methods cannot be extend to graph-level tasks, we do not report their results. Based on the results presented in Table 3, we summarize the following key findings: (1) All HNN methods achieve a substantial performance advantage on synthetic hypergraph classification datasets over real-world ones. The inferior performance on real-world datasets reflects the higher structural complexity of real-world hypergraphs, underscoring the importance of designing and evaluating algorithms on real-world data. (2) The classification performance of HNN algorithms varies greatly across datasets, with no method consistently achieving state-of-the-art results. For example, while DPHGNN achieves the highest accuracy on IMDB-Dir-Form, it ranks only 11th among 14 methods on IMDB-Dir-Genre, suggesting that dataset characteristics can significantly affect model performance. (3) Higher accuracy does not necessarily guarantee improved macro-F1 performance, particularly in class-imbalanced settings. For example, on the imbalanced Twitter-Friend dataset, HNHN achieves the third-highest accuracy among all HNN methods, yet obtains a macro-F1 of only 39.40%, which is the lowest among all methods. Conversely, HyperND attains the best macro-F1 on the same dataset but ranks only 10th in accuracy out of the 14 evaluated models.
| Method | Cora | PubMed | Cora-CA | DBLP-CA | Actor | Pokec | ||||||
| AUROC | AP | AUROC | AP | AUROC | AP | AUROC | AP | AUROC | AP | AUROC | AP | |
| HGNN | 73.701.19 | 71.731.57 | 66.089.84 | 63.679.02 | 89.161.11 | 89.850.82 | 75.443.01 | 73.964.91 | 72.421.96 | 68.791.83 | 86.090.92 | 84.320.95 |
| HyperGCN | 77.341.30 | 77.150.33 | 66.468.87 | 64.847.98 | 92.730.95 | 93.420.89 | 89.460.18 | 91.390.34 | 55.018.76 | 56.297.44 | 91.450.70 | 90.760.70 |
| HCHA | 73.571.08 | 72.241.80 | 63.351.61 | 63.131.47 | 85.853.27 | 84.775.66 | 73.303.72 | 72.093.07 | 69.860.95 | 66.720.74 | 88.810.28 | 88.250.41 |
| LEGCN | 67.162.85 | 68.764.89 | 56.393.28 | 54.332.41 | 74.290.59 | 75.950.62 | 50.701.40 | 50.470.94 | 48.253.00 | 49.761.29 | 74.941.44 | 73.891.04 |
| HyperND | 69.101.28 | 72.711.48 | 72.120.78 | 73.530.63 | 84.010.61 | 84.981.07 | 78.630.71 | 79.420.94 | 53.122.56 | 52.642.33 | 75.771.56 | 73.511.62 |
| PhenomNN | 75.710.91 | 75.221.42 | 74.290.85 | 72.931.27 | 80.271.62 | 79.591.11 | 75.860.86 | 75.540.88 | 56.653.04 | 55.752.87 | 70.832.52 | 70.172.36 |
| SheafHyperGNN | 70.535.28 | 70.93 4.04 | 68.261.92 | 68.071.18 | 79.214.53 | 75.426.73 | 76.301.91 | 75.411.76 | 59.836.77 | 59.845.73 | 83.442.49 | 85.111.80 |
| HJRL | 58.482.52 | 61.022.60 | 59.280.84 | 58.631.50 | 82.411.90 | 85.671.11 | OOM | OOM | 48.260.77 | 50.000.31 | 84.883.30 | 86.182.61 |
| DPHGNN | 66.485.82 | 67.235.11 | 60.377.77 | 59.867.70 | 82.892.28 | 83.782.50 | OOM | OOM | 42.445.81 | 46.603.03 | 73.354.59 | 73.283.74 |
| TF-HNN | 76.940.86 | 76.570.71 | 73.750.73 | 75.540.72 | 74.971.85 | 71.131.65 | 75.702.77 | 74.692.26 | 54.031.71 | 54.061.57 | 68.000.97 | 67.411.20 |
| HNHN | 70.131.67 | 68.841.09 | 55.670.39 | 53.520.31 | 84.331.40 | 83.491.00 | 82.850.70 | 82.130.58 | 69.890.98 | 66.450.74 | 82.251.34 | 81.721.53 |
| UniGNN | 73.510.87 | 75.231.51 | 74.200.82 | 71.761.16 | 80.590.98 | 82.371.11 | 81.080.79 | 79.390.46 | 50.241.26 | 50.010.56 | 85.641.20 | 84.361.48 |
| AllSetTransformer | 72.552.95 | 74.861.85 | 71.092.99 | 73.152.49 | 76.137.70 | 75.028.68 | 75.124.14 | 77.124.22 | 55.845.99 | 58.734.39 | 83.654.34 | 83.364.72 |
| ED-HNN | 67.241.91 | 69.892.24 | 70.090.43 | 72.610.48 | 74.581.37 | 72.941.32 | 81.860.67 | 84.750.50 | 51.742.79 | 52.272.54 | 85.271.48 | 84.951.43 |
| EHNN | 78.990.99 | 79.540.93 | 76.500.62 | 75.940.70 | 77.833.01 | 78.293.72 | 87.960.98 | 89.000.64 | 65.690.46 | 65.370.35 | 88.631.58 | 91.310.88 |
| T-HyperGNN | 58.911.23 | 62.171.58 | 58.354.43 | 55.813.71 | 66.870.88 | 69.650.53 | 67.175.79 | 68.453.85 | 49.160.22 | 50.200.41 | 65.211.21 | 66.901.56 |
| Method | RHG-10 | RHG-3 | IMDB-Dir-Form | IMDB-Dir-Genre | Steam-Player | Twitter-Friend | ||||||
| Acc | F1_ma | Acc | F1_ma | Acc | F1_ma | Acc | F1_ma | Acc | F1_ma | Acc | F1_ma | |
| HGNN | 94.601.66 | 94.471.84 | 98.930.68 | 98.970.65 | 63.720.62 | 57.922.24 | 76.762.66 | 72.024.37 | 51.652.51 | 50.912.92 | 55.422.03 | 46.814.27 |
| HyperGCN | 85.501.10 | 95.421.09 | 99.470.50 | 99.480.49 | 62.870.40 | 57.202.46 | 77.530.99 | 72.971.08 | 51.173.32 | 50.483.12 | 56.954.17 | 50.125.88 |
| HCHA | 96.601.02 | 96.481.09 | 99.330.42 | 99.370.38 | 61.602.16 | 55.372.17 | 78.121.96 | 73.203.00 | 52.432.30 | 51.772.52 | 58.172.34 | 49.576.84 |
| LEGCN | 92.401.16 | 92.061.19 | 96.800.98 | 96.780.29 | 61.811.32 | 56.053.75 | 76.381.68 | 72.031.54 | 53.111.58 | 52.701.87 | 56.643.72 | 53.384.92 |
| HyperND | 91.000.95 | 90.741.04 | 92.801.95 | 92.751.90 | 60.743.25 | 55.024.95 | 75.650.51 | 71.371.10 | 53.882.15 | 49.712.05 | 55.273.79 | 43.616.51 |
| PhenomNN | 91.100.73 | 90.770.77 | 93.471.90 | 93.451.90 | 61.281.97 | 53.713.13 | 74.590.61 | 70.150.88 | 51.653.06 | 48.944.55 | 57.403.84 | 48.264.66 |
| SheafHyperGNN | 96.001.38 | 95.961.32 | 99.730.33 | 99.740.32 | 62.342.06 | 56.473.49 | 77.001.14 | 72.781.17 | 53.112.39 | 52.562.74 | 56.492.51 | 51.434.42 |
| HJRL | 96.100.80 | 95.980.85 | 99.600.53 | 99.600.53 | 63.092.83 | 56.543.62 | 77.821.47 | 73.731.92 | 51.843.52 | 51.133.29 | 57.102.79 | 44.197.19 |
| DPHGNN | 96.800.68 | 96.710.71 | 99.490.65 | 99.610.64 | 64.042.70 | 57.413.96 | 76.181.30 | 71.591.82 | 51.361.72 | 49.033.63 | 59.242.88 | 46.128.49 |
| TF-HNN | 95.900.80 | 95.880.78 | 98.800.65 | 98.840.61 | 62.341.76 | 55.323.81 | 76.411.31 | 71.891.45 | 54.851.82 | 52.722.54 | 56.183.53 | 44.178.95 |
| HNHN | 94.001.90 | 94.081.88 | 99.920.02 | 99.950.02 | 62.342.98 | 55.243.88 | 73.651.47 | 69.681.18 | 52.821.61 | 52.681.69 | 58.474.65 | 39.403.14 |
| UniGNN | 95.501.38 | 95.401.44 | 98.800.27 | 98.830.27 | 61.062.88 | 55.754.01 | 77.120.88 | 72.931.43 | 51.462.48 | 48.852.59 | 55.884.14 | 46.484.90 |
| AllSetTransformer | 97.300.98 | 97.261.04 | 98.800.27 | 98.810.26 | 62.231.01 | 56.262.93 | 76.471.38 | 72.261.12 | 53.012.77 | 48.217.20 | 60.151.70 | 51.527.00 |
| ED-HNN | 96.500.77 | 96.410.78 | 99.070.53 | 99.100.51 | 62.132.36 | 57.004.71 | 77.121.11 | 72.870.44 | 52.822.65 | 48.732.36 | 57.402.66 | 42.575.09 |
4.2 Efficiency and Scalability Evaluation (RQ2)
To answer RQ2, we analyze the training time and memory consumption of all methods on datasets of varying scales, including Cora, DBLP-CA, Yelp, and Trivago. From Figure 2, we make the following observations: (1) Most HNN methods struggle to achieve a satisfactory balance between predictive performance and efficiency. For example, on the Yelp dataset, although EDHNN and EHNN achieve higher accuracy than the simple HGNN method, their training times are more than 9x and 23x longer, respectively. Similarly, many HNNs consume excessive memory, particularly on large-scale datasets. On Yelp, 7 out of 16 methods encounter out-of-memory (OOM) issues. On the Trivago dataset, although several HNNs remain computationally scalable, most methods fail to deliver satisfactory predictive performance. Only 2 out of 10 methods, TFHNN (90.79%) and HyperND (87.19%), achieve accuracy above 60%. (2) Compared to spectral-based and spatial-based methods, tensor-based approaches generally suffer from more severe efficiency and scalability challenges. T-HyperGNN can only scale to the medium-sized DBLP-CA dataset, where it runs approximately 406 times slower than the fastest HGNN method. Likewise, EHNN is the slowest method on Yelp and fails to scale to the large-scale Trivago dataset. (3) Among all evaluated methods, TFHNN demonstrates the most favorable trade-off between effectiveness, efficiency, and scalability. In particular, it achieves the highest accuracy on three out of the four datasets. On the large-scale Trivago dataset, it delivers the best predictive performance while consuming no more than 1.6 GB of memory and requiring less than 30 seconds of runtime, ranking first in memory usage and second in training time among all methods. These results highlight the strong potential of the decoupled architecture for designing efficient and scalable HNN algorithms. Building on these findings, future research should place greater emphasis on innovative architectural design to achieve a better balance among effectiveness, efficiency, and scalability, thereby facilitating the development of more practical HNN solutions.
4.3 Robustness Evaluation (RQ3)
To investigate the robustness of HNN algorithms, we focus on three key aspects: structural robustness, feature robustness, and supervision signal robustness. While our experiments primarily focus on the node classification task due to space constraints, the DHG-Bench library supports flexible extension to other tasks. We include 8 representative models from the three HNN categories.
4.3.1 Robustness analysis with respect to structure perturbations. To analyze structure-level robustness, we randomly remove or add node-hyperedge incidence connections in the original hypergraph structures of Cora and Actor, and evaluate the performance of HNN algorithms on the perturbed hypergraphs. The modification ratio is varied from 0 to 0.9 to simulate different levels of attack intensity. From Figure 3, we find that: (1) As the ratio of deletions or additions increases, model performance generally declines; however, the extent of degradation varies significantly across datasets. In particular, the accuracy drop caused by structural perturbations is more pronounced on the homophilic hypergraph Cora than on the heterophilic hypergraph Actor. (2) Spectral-based methods tend to be more susceptible to structural noise. For example, on the Cora dataset, as the deletion ratio increases from 0 to 0.9, spectral-based methods such as HGNN, SheafHyperGNN, and DPHGNN suffer an accuracy drop of over 30%, while methods from the other two categories consistently experience no more than a 10% decline.
4.3.2 Robustness analysis with respect to feature perturbations. To study feature-level robustness, we consider feature noise and sparsity. For feature noise, Gaussian noise is injected based on the dimensionality of node features. Sparsity is introduced by randomly masking a certain proportion of node features with zeros. As shown in Figure 4, we can observe that: (1) Although both increased feature noise and greater sparsity lead to performance degradation, sparsity has a more pronounced negative impact on accuracy than random feature noise. Specifically, on the Actor dataset, when the sparsity ratio increases from 0 to 0.9, all methods suffer an accuracy drop of about 10%. In contrast, under the same increase in noise level, 7 out of 8 methods experience accuracy losses of less than 5%. (2) The robustness of most algorithms to feature sparsity varies significantly across different datasets. In particular, most HNNs show better robustness on Actor than on Cora. At a feature mask ratio of 0.9, all algorithms on Cora suffer a performance drop exceeding 20%, whereas on Actor, more than half of the models show a decrease of less than 13%.
4.3.3 Robustness analysis with respect to supervision signals. On the basis of exploring feature and structural robustness, we further study supervision signal robustness by simulating realistic noise and sparsity scenarios. For label noise, a certain proportion of training samples are randomly assigned incorrect labels, while sparsity is introduced by varying the proportion of training nodes. Figure 5 indicates several key observations: (1) Model performance consistently declines as noise intensity increases or supervision signals become sparser, demonstrating a clear downward trend. Specifically, across different datasets, label noise leads to an approximately linear decline in accuracy as the noise ratio increases. In contrast, the effect of supervision sparsity is relatively modest at low sparsity levels but becomes markedly more severe at higher ratios, resulting in accelerated performance degradation. (2) Label noise has a more pronounced impact than feature-level or structural-level noise. Under the same noise level, label noise consistently causes greater performance degradation compared to feature or structure perturbations. For instance, when the label noise ratio reaches 15%, all HNN methods suffer an accuracy drop exceeding 15%. In contrast, most methods do not experience comparable degradation even when the feature or structural noise reaches 90%.
4.4 Fairness Evaluation (RQ4)
| Method | German | Bail | Credit | ||||||
| Acc | Acc | Acc | |||||||
| MLP | 67.683.46 | 1.781.30 | 2.590.99 | 89.401.76 | 6.160.93 | 1.790.84 | 79.690.85 | 3.430.83 | 2.070.43 |
| HGNN | 69.762.50 | 9.593.51 | 6.903.82 | 91.020.54 | 7.830.80 | 2.601.05 | 80.210.41 | 5.042.07 | 3.461.05 |
| HyperGCN | 70.403.23 | 6.392.87 | 3.571.61 | 94.720.79 | 7.900.95 | 1.230.52 | 80.420.34 | 5.382.61 | 3.891.60 |
| HCHA | 70.562.51 | 9.373.74 | 6.723.05 | 91.520.92 | 7.620.95 | 1.400.80 | 80.080.43 | 3.581.87 | 2.470.78 |
| LEGCN | 70.883.22 | 7.532.54 | 3.252.00 | 95.020.40 | 7.870.62 | 1.280.52 | 80.480.37 | 4.312.24 | 3.151.12 |
| HyperND | 71.042.61 | 7.374.70 | 3.673.10 | 89.752.41 | 7.921.52 | 3.192.22 | 80.020.49 | 4.142.22 | 2.500.72 |
| PhenomNN | 70.962.85 | 3.543.07 | 1.601.94 | 91.711.13 | 10.831.64 | 1.940.40 | OOM | OOM | OOM |
| SheafHyperGNN | 70.643.29 | 8.143.25 | 5.232.05 | OOM | OOM | OOM | OOM | OOM | OOM |
| HJRL | 69.923.46 | 3.522.70 | 3.051.82 | OOM | OOM | OOM | OOM | OOM | OOM |
| TF-HNN | 70.483.14 | 5.273.09 | 4.192.23 | 95.330.25 | 7.960.65 | 1.030.67 | 80.460.36 | 4.932.44 | 3.431.43 |
| DPHGNN | 70.243.25 | 2.250.49 | 1.380.77 | 93.410.93 | 8.071.20 | 2.051.22 | OOM | OOM | OOM |
| HNHN | 69.523.62 | 4.012.76 | 1.591.60 | 90.761.30 | 6.031.43 | 3.041.34 | 78.000.23 | 5.703.11 | 4.673.10 |
| UniGNN | 71.072.70 | 5.083.03 | 2.801.32 | 91.301.47 | 9.421.68 | 3.942.18 | 80.440.37 | 3.902.40 | 2.851.33 |
| AllSetTransformer | 70.483.11 | 4.473.39 | 3.503.38 | 96.261.83 | 8.360.85 | 1.951.10 | 80.400.44 | 4.462.96 | 3.441.60 |
| ED-HNN | 70.163.15 | 4.063.05 | 4.072.75 | 94.260.77 | 8.080.64 | 1.530.26 | OOM | OOM | OOM |
| EHNN | 70.403.07 | 2.875.73 | 2.344.69 | 93.621.75 | 9.291.60 | 2.881.23 | 80.340.47 | 4.512.77 | 3.131.75 |
| T-HyperGNN | 71.201.82 | 8.996.52 | 6.805.02 | OOM | OOM | OOM | OOM | OOM | OOM |
We evaluate the fairness of various HNN algorithms on three datasets containing sensitive attributes. Predictive performance is measured by accuracy, while fairness is assessed using demographic parity () and equalized odds (). The experimental results in Table 4 show that: (1) While most existing HNN algorithms demonstrate stronger predictive performance, they tend to show more significant fairness concerns than traditional MLPs that do not rely on message passing. For instance, on the Credit dataset, MLP achieves of 3.43 and of 2.07, both lower than those of UniGNN, the fairest HNN model. This suggests that hypergraph message passing may exacerbate fairness issues in high-order representation learning. (2) Algorithmic fairness varies significantly across different datasets and evaluation metrics. No HNN method consistently demonstrates superior fairness across different settings. For instance, while HCHA achieves the best fairness performance on the Credit dataset across both metrics, its and rank as the second and third worst, respectively, on the German dataset. Moreover, the results reveal that and capture different aspects of fairness, and better performance in one metric does not necessarily indicate better performance in the other. For example, on the Bail dataset, although HNHN attains the lowest among all HNN methods, its ranks as the third highest among the 16 models. These observations highlight the necessity of developing fairness-aware HNN algorithms that can achieve robust fairness across diverse datasets and evaluation metrics.
5 Conclusion and Future Directions
This paper presents DHG-Bench, the first comprehensive benchmark for deep hypergraph learning, which integrates and compares 16 representative HNNs across 20 diverse hypergraph datasets encompassing various domains, sizes, and structural properties. We conduct extensive experiments to evaluate the effectiveness, efficiency, robustness, and fairness of DHGL algorithms on node-level, edge-level, and graph-level tasks. Our analysis reveals the strengths and weaknesses of different HNN models under a wide range of settings, offering valuable insights into their practical applicability and design trade-offs. Furthermore, we develop and release a package DHG-Bench (https://github.com/Coco-Hut/DHG-Bench) that includes all experimental protocols, baseline algorithms, datasets, and reproducibility scripts to facilitate future research.
Drawing upon our empirical analyses, we point out some promising future directions for DHGL community.
Designing a unified algorithm for diverse tasks and datasets. Our experiments reveal that existing HNN architectures show substantial performance disparities across datasets and tasks, which hinders their transferability to diverse scenarios. Future research should explore novel architectural designs and training strategies that promote robust generalization across heterogeneous datasets and tasks, ultimately moving toward a versatile “one-for-all” solution in DHGL.
Improving the Efficiency and Scalability of HNN methods. Our results show that many HNN algorithms either fail to scale to large datasets or suffer significant performance degradation when applied to them. Given that real-world hypergraphs in industrial applications often contain millions of nodes and hyperedges, future work should focus on designing HNN algorithms that deliver both high efficiency and strong effectiveness at scale. Progress in this direction will be essential for enabling the practical adoption of HNNs in industrial settings.
Enhancing the robustness of HNN algorithms. Our experimental results show that current HNN methods are susceptible to data perturbations across multiple aspects, including structural modifications, feature noise, and corrupted supervisory signals. Future research should focus on designing robust HNN algorithms that are resilient to such perturbations and maintain reliable performance in noisy industrial environments.
Developing fairness-aware HNN algorithms. Empirical evidence suggests that current message passing-based HNN algorithms often suffer from fairness issues, leading to biased predictions that disadvantage certain demographic groups or individuals. Future research should explore how to develop fairness-aware HNN methods that mitigate such biases and discrimination, enabling their safe adoption in high-stakes real-world applications such as crime prediction (Suresh & Guttag, 2019), job hunting (Mehrabi et al., 2021), and credit evaluation (Yeh & Lien, 2009).
Appendix A Datasets and Algorithms
A.1 Detailed description of datasets
We adopt 20 publicly available benchmark datasets, ranging from node-level to graph-level, to comprehensively evaluate HNN algorithms in terms of effectiveness, efficiency, robustness, and fairness. Detailed descriptions of these datasets are provided below.
Node/Edge-level Datasets
Cora/Pubmed/Cora-CA/DBLP-CA (Yadati et al., 2019): Cora and Pubmed are co-citation networks where nodes represent papers and hyperedges connect papers cited together. Cora-CA and DBLP-CA are co-authorship hypergraphs, with nodes as papers and hyperedges linking all papers co-authored by the same author. Node features are Bag-of-Words representations of the documents, and labels indicate paper categories.
Warmart (Chien et al., 2022): The Walmart dataset models a hypergraph where nodes represent products and hyperedges capture sets of products purchased together. Node labels indicate product categories. Following (Chien et al., 2022), each node feature is a 100-dimensional vector obtained by adding Gaussian noise with to one-hot encodings of the labels.
Trivago (Kim et al., 2023): Trivago is a hotel-web search hypergraph where each node indicates each hotel and each hyperedge corresponds to a user. If a user (hyperedge) has visited the website of a particular hotel (node), the corresponding node is added to the respective user hyperedge. Furthermore, each hotel’s class is labeled based on the country in which it is located.
Actor (Li et al., 2025): The actor co-occurrence network is derived from a heterogeneous movie-actor-director-writer network 222https://www.aminer.org/lab-datasets/soinf/, capturing intricate collaborations within films. Nodes represent individuals involved in film production (actors, directors, writers), and hyperedges denote their joint participation in a single film. Node attributes are extracted from Wikipedia keywords, and labels indicate each individual’s specific role.
Amazon-ratings (Ratings) (Li et al., 2025): This dataset, sourced from the Amazon co-purchasing network in the SNAP repository (Leskovec, 2014), includes products like books, music CDs, DVDs, and VHS tapes. Nodes represent individual products, and hyperedges link those frequently co-purchased together. The task is to predict each product’s average user rating, classified into ten levels. Node features are extracted using the Bag of Words method applied to product descriptions (Juluru et al., 2021).
Twitch-gamers (Gamers) (Li et al., 2025): The Twitch-gamers dataset is a connected undirected hypergraph representing user interactions on the Twitch streaming platform. Nodes denote user accounts, and hyperedges are formed based on mutual follows within specific timeframes. Each node is associated with features such as view counts, timestamps, language preferences, activity duration, and inactivity status. The goal is to predict whether a channel hosts explicit content (binary classification).
Pokec (Li et al., 2025): The Pokec dataset is derived from Slovakia’s largest online social networking platform and is used to model social relationships and attributes. Nodes represent individual users, and hyperedges correspond to each user’s full set of friends. Node labels indicate user-reported gender, while node features are extracted from profile information, including age, hobbies, interests, education level, region, etc.
Yelp (Chien et al., 2022): The Yelp dataset is a hypergraph where nodes represent restaurants and hyperedges link those visited by the same user. Node labels denote average star ratings (1.0–5.0 in 0.5 steps). Features include geographic coordinates, one-hot encodings of city/state, and Bag-of-Words vectors from the top-1000 restaurant name tokens.
German Credit (German) (Wei et al., 2022): The nodes in the dataset represent clients in German Bank, and hyperedges are constructed by linking individuals with the most similar credit accounts to each person in the dataset. The task is to classify credit risk levels as high or low based on the sensitive attribute ”gender” (Male/Female).
Bail (Wei et al., 2022): The nodes in the datasets are defendants who got released on bail at the U.S state courts during 1990- 2009 (Jordan & Freiburger, 2015). Hyperedges are constructed based on the similarity of past criminal records among individuals. The task is to classify whether defendants are on bail or not with the sensitive attribute ”race” (White/Black).
Credit Defaulter (Credit) (Wei et al., 2022): The nodes in the dataset represent credit card users, and hyperedges are formed based on the similarity of users’ spending and payment patterns. The task is to classify the default status with the sensitive attribute ”age” (¡25/¿25).
Graph-level Datasets
RHG-10/RHG-3 (Feng et al., 2024): RHG-10 dataset encompasses ten distinct synthetic factor hypergraph structures (i.e., Hyper Flower, Hyper Pyramid, Hyper Checked Table, Hyper Wheel, Hyper Lattice, Hyper Windmill, Hyper Firm Pyramid, Hyper RChecked Table, Hyper Cycle, and Hyper Fern). To evaluate the algorithm’s ability to recognize significant high-order structures, RHG-3 dataset is constructed by randomly generating hypergraphs for three distinctively various hypergraph structures: hyper-pyramid, hyper-check-table, and hyper-wheel.
IMDB-Dir-Form/IMDB-Dir-Genre (Feng et al., 2024): These two datasets contain hypergraphs constructed by the co-director relationship from the original IMDB dataset. The director of each movie is a hypergraph. ”Form” included in the dataset’s name indicates that the movie category is identified by its form, like animation, documentary, and drama. ”Genre” denotes the movie is classified by its genres, like adventure, crime, and family.
Stream-Player (Feng et al., 2024): The Steam-Player dataset is a player dataset where each player is a hypergraph. The vertex is the games played by the player, and the hyperedge is constructed by linking the games with shared tags. The target of the dataset is to identify each user’s preference: single-player-game or multi-player-game.
Twitter-Friend (Feng et al., 2024): The Twitter-Friend dataset is a social media dataset. Each hypergraph is the friends of a specified user. The hyperedge is constructed by linking the users who are friends. The label associated with the hypergraph is to identify whether the user posted the blog about ”National Dog Day” or ”Respect Tyler Joseph”.
| Datasets | Nodes | Edges | Features | Classes | Type | Description |
| Cora | 2,708 | 1,579 | 1,433 | 7 | Homophily | Co-citation |
| PubMed | 19,717 | 7,963 | 500 | 3 | Homophily | Co-citation |
| Cora-CA | 2,708 | 1,072 | 1,433 | 7 | Homophily | Co-author |
| DBLP-CA | 41,302 | 22,363 | 1,425 | 6 | Homophily | Co-author |
| Warmart | 88,860 | 69,906 | 100 | 11 | Homophily | Co-purchasing |
| Trivago | 172,738 | 233,202 | 300 | 160 | Homophily | Co-click |
| Actor | 16,255 | 10,164 | 50 | 3 | Heterophily | Co-occurence |
| Ratings | 22,299 | 2,090 | 111 | 5 | Heterophily | Co-purchasing |
| Gamers | 16,812 | 2,627 | 7 | 2 | Heterophily | Co-create |
| Pokec | 14,998 | 2,406 | 65 | 2 | Heterophily | Co-friendship |
| Yelp | 50,758 | 679,302 | 1,862 | 9 | Heterophily | Co-visit |
| Datasets | German Credit | Bail | Credit Defaulter |
| Nodes | 1000 | 18876 | 30000 |
| Edges | 1000 | 18876 | 30000 |
| Features | 27 | 18 | 13 |
| Sens Attribute | Gender | Race | Age |
| Labels | Credit status | Bail decision | Future default |
| Datasets | Hypergraphs | Avg. | Avg. | Avg. | Classes |
| RHG-10 | 2,000 | 31.3 | 29.8 | 5.2 | 10 |
| RHG-3 | 1,500 | 35.5 | 17.9 | 6.9 | 3 |
| IMDB-Dir-Form | 1,869 | 15.7 | 39.2 | 3.7 | 3 |
| IMDB-Dir-Genre | 3,393 | 17.3 | 36.4 | 3.8 | 3 |
| Stream-Player | 2,048 | 13.8 | 46.4 | 4.5 | 2 |
| Twitter-Friend | 1,310 | 21.6 | 84.3 | 4.3 | 2 |
A.2 Detailed description of algorithms
In our developed DHG-Bench, we integrate 16 state-of-the-art HNN algorithms, including 10 spectral-based methods, 4 spatial-based methods, and 2 tensor-based methods. Specifically, we introduce all the HNN models as follows.
Spectral-based HNN Models
HGNN (Feng et al., 2019): HGNN is a framework for representation learning that extends spectral convolution to hypergraphs. By leveraging the hypergraph Laplacian and approximating spectral filters with truncated Chebyshev polynomials, it effectively captures high-order correlations inherent in complex data.
HyperGCN (Yadati et al., 2019): HyperGCN approximates each hyperedge of the hypergraph by a set of pairwise edges connecting the vertices of the hyperedge and treats the learning problem as a graph learning problem on the approximation.
HCHA (Bai et al., 2021): HCHA is a hypergraph neural network that introduces two end-to-end trainable operators: hypergraph convolution and hypergraph attention. Hypergraph convolution efficiently propagates information by leveraging high-order relationships and local clustering structures, with standard graph convolution shown as a special case. Hypergraph attention further enhances representation learning by dynamically adjusting hyperedge connections through an attention mechanism, enabling task-relevant information aggregation and yielding more discriminative node embeddings.
LEGCN (Yang et al., 2022): LEGCN is a hypergraph learning model based on the Line Expansion (LE). By modeling vertex-hyperedge pairs, LEGCN bijectively transforms a hypergraph into a simple graph, preserving the symmetric co-occurrence structure and avoiding information loss. This enables existing graph learning algorithms to operate directly on hypergraphs.
HyperND (Prokopchik et al., 2022): HyperND develops a nonlinear diffusion process on hypergraphs that spreads both features and labels following the hypergraph structure. The novel diffusion can incorporate a broad class of nonlinearities to increase the modeling capability, and the limiting point serves as a node embedding from which we make predictions with a linear model.
PhenomNN (Wang et al., 2023b): PhenomNN is a hypergraph learning framework grounded in a family of expressive, parameterized hypergraph-regularized energy functions. It formulates node embeddings as the minimizers of these energy functions, which are optimized jointly with a parameterized classifier through a supervised bilevel optimization process. This approach provides a principled way to model high-order relationships in hypergraphs while enabling end-to-end training.
SheafHyperGNN (Duta et al., 2023): SheafHyperGNN introduces a cellular sheaf framework for hypergraphs, enabling the modeling of complex dynamics while preserving their higher-order connectivity. Then, it generalize the two commonly used hypergraph Laplacians to incorporate the richer structure sheaves offer and construct two powerful neural networks capable of inferring and processing hypergraph sheaf structure.
HJRL (Yan et al., 2024): HJRL introduces a novel cross expansion method, which transforms both hypervertices and edges of a hypergraph to vertices in a standard graph. Then, a joint learning model is proposed to embed both hypervertices and hyperedges into a shared representation space. In addition, the algorithm employs a hypergraph reconstruction objective to preserve structure information in the model.
TF-HNN (Tang et al., 2025): TF-HNN is the first model to decouple hypergraph structural processing from model training, substantially improving training efficiency. Specifically, it introduces a unified, training-free message-passing module (TF-MP-Module) by identifying feature aggregation as the core operation in HNNs. The TF-MP-Module removes learnable parameters and nonlinear activations, and compresses multi-layer propagation into a single step, offering a simplified and efficient alternative to existing architectures.
DPHGNN (Saxena et al., 2024): DPHGNN is a hybrid spectral-spatial framework designed for effective feature representation in resource-constrained hypergraph settings. It introduces equivariant operator learning to capture lower-order semantics by inducing topology-aware inductive biases. It employs a dual-layered feature update mechanism: a static update layer provides spectral biases and relational features, while a dynamic update layer fuses explicitly aggregated features from the underlying graph topology into the hypergraph message-passing process.
Spatial-based HNN Models
HNHN (Dong et al., 2020): HNHN is a hypergraph convolution network with nonlinear activation functions applied to both hypernodes and hyperedges, combined with a normalization scheme that can flexibly adjust the importance of high cardinality hyperedges and high-degree vertices depending on the dataset.
UniGNN (Huang & Yang, 2021): UniGNN is a unified message-passing framework that generalizes standard GNNs to hypergraphs. It models the two-stage aggregation process by first computing hyperedge representations using a permutation-invariant function over the features of incident vertices, and then updating each vertex by aggregating its associated hyperedge representations. This formulation enables seamless adaptation of existing GNN architectures to hypergraph structures.
AllSetTransformer (Chien et al., 2022): AllSet is a novel HNN paradigm that implements HNN layers as compositions of two multiset functions. The AllSetTransformer integrates powerful Set Transformer (Lee et al., 2019) in the design of the specialized HNN, enabling greater modeling flexibility and enhanced expressive power.
ED-HNN (Wang et al., 2023a): ED-HNN is a HNN architecture designed to approximate any continuous, permutation-equivariant hypergraph diffusion operator. The model is efficiently implemented by combining the star expansion (bipartite representation) of hypergraphs with standard message passing neural networks, and supports scalable training via shared weights across layers.
Tensor-based HNN Models
EHNN (Kim et al., 2022): EHNN is the first framework to realize equivariant GNNs for general hypergraph learning. It establishes a connection between sparse hypergraphs and dense, fixed-order tensors, enabling the design of a maximally expressive equivariant linear layer. To ensure scalability and generalization to arbitrary hyperedge orders, EHNN further introduces hypernetwork-based parameter sharing.
T-HyperGNN (Wang et al., 2024a): T-HyperGNN is a general framework that integrates tensor hypergraph signal processing (t-HGSP) (Pena-Pena et al., 2023) to encode hypergraph structures using tensors. It models node interactions through multiplicative interaction tensors, elevating aggregation from traditional linear operations to higher-order polynomial mappings, thereby enhancing expressive power. To ensure scalability, T-HyperGNN introduces tensor-message-passing by exploiting tensor sparsity, enabling efficient processing of large hypergraphs with computational and memory costs comparable to matrix-based HNNs.
Appendix B Additional Experimental Details
B.1 General experimental settings
| Method | Hyperparameter | Search Space |
| General Settings | Epochs | 100, 200, 300, 400, 500, 800, 1000 |
| Learning Rate | 0.1, 0.01, 0.001, 0.0001 | |
| Layers | 1, 2, 3, 4 | |
| Dropout Rate | 0, 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8 | |
| Weight Decay | 0, 0.0005 | |
| Hidden Units | 64, 128, 256, 512, 1024 | |
| Activation | LeakyReLU, ReLU, PReLU, Sigmoid, Softmax | |
| Hyperedge Pooling | max, mean, max-min | |
| Hypergraph Pooling | max, mean | |
| HCHA | heads | 1, 2, 4, 8, 16 |
| HyperND | HyperND_ord | 1, 2, 3, 5, 10 |
| HyperND_tol | 0.001, 0.0001, 0.00001, 0.000001 | |
| HyperND_steps | 50, 100, 150, 200 | |
| alpha | 0, 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9 | |
| HJRL | 0.001, 0.01, 0.1, 1, 10 | |
| PhenomNN | 0, 0.1, 1, 10, 20, 50, 80, 100 | |
| 0, 0.1, 1, 10, 20, 50, 80, 100 | ||
| prop_steps | 2, 4, 8, 16, 32, 64, 128 | |
| alpha | 0, 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9 | |
| SheafHyperGNN | init_hedge | ’rand’, ’avg’ |
| sheaf_pred_block | ’MLP_var1’, ’MLP_var2’, ’MLP_var3’, ’cp_decomp’ | |
| sheaf_transformer_head | 1, 2, 4, 8, 16 | |
| stalk_dim | 1, 2, 4, 8 | |
| TFHNN | mlp_hidden_size | 64, 128, 256, 512, 1024 |
| mlp_num_layers | 1, 2, 3, 4 | |
| HNHN | alpha | -3.0, -2.5, -2.0, -1.5, -1.0, -0.5, 0.0, 0.5 |
| beta | -2.5, -2.0, -1.5, -1.0, -0.5, 0.0, 0.5, 1.0 | |
| UniGNN | alpha | 0, 0.1, 0.2, 0.3, 0.4, 0.5, 0.8, 0.9 |
| beta | 0, 0.1, 0.2, 0.3, 0.4, 0.5, 0.8, 0.9 | |
| AllSetTransformer | attention_heads | 1, 2, 4, 8, 16 |
| EDHNN | alpha | 0, 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9 |
| DPHGNN | attention_heads | 1, 2, 4, 8, 16 |
| EHNN | ehnn_qk_channels | 64,128,256 |
| ehnn_n_heads | 1,2,4,8,16 | |
| ehnn_pe_dim | 64, 128 | |
| ehnn_inner_channel | 64, 128, 256 | |
| ehnn_hidden_channel | 64, 128, 256 | |
| T-HyperGNN | M: maximum cardinality | 1,2,3,4,5 |
| combine | concat, sum |
We strive to follow the original implementations of various HNN methods from their respective papers or source codes and integrate them into a unified training and evaluation framework. All parameters are randomly initialized. We use the cross-entropy loss function (Mao et al., 2023) for all three benchmark classification tasks. Adam optimizer (Kingma, 2014) is adopted with an appropriate learning rate and weight decay to achieve the best performance on the validation split. We repeat all experiments five times and report the average results along with standard deviations.
B.2 Evaluation metrics
We incorporate a variety of metrics to ensure a comprehensive benchmark evaluation across different tasks.
For node classification task, we adopt widely-used Accuracy (Acc) (Feng et al., 2019; Chien et al., 2022; Wang et al., 2025) to measure the predictive capacity of each HNN algorithm. As for fairness evaluation, we utilize two widely used group fairness metrics: demographic parity () (Dwork et al., 2012), and equalized odds () (Hardt et al., 2016). Specifically, is defined as the acceptance rate difference between the two sensitive groups, which can be formulated as:
where denotes predicted node label and is the sensitive attribute. Similarly, quantifies the disparity in predictions by further conditioning on the ground truth , formulated as:
Both metrics are estimated on the test dataset. Lower values of and indicate better fairness.
For hyperedge prediction task, following (Hwang et al., 2022; Ko et al., 2025; Yu et al., 2025), we use AUROC (Area Under the Receiver Operating Characteristic Curve) and AP (Average Precision), where higher values of these metrics indicate stronger hyperedge prediction performance.
For the hypergraph classification task, following (Feng et al., 2024), we report both accuracy and macro-F1 score, where macro-F1 provides a balanced assessment of classification performance across all classes, especially in the presence of class imbalance.
B.3 Hyperparameter setting
We carefully tune hyperparameters to ensure a rigorous and unbiased evaluation of the integrated HNN methods. For algorithms without explicit hyperparameter guidelines in their original papers or source code, we perform a grid search to identify optimal configurations. The search spaces are provided in Table A4. For detailed interpretations, please refer to the corresponding papers. All complete hyperparameter configurations are available in our publicly released GitHub repository.
B.4 Experimental environment
All the experiments are conducted with the following computational resources and configurations:
Appendix C Reproducibility
Accessibility. All the datasets, algorithm implementations, experimental settings, and hyper-parameter configurations are publicly available in our open-source library DHG-Bench at https://github.com/Coco-Hut/DHG-Bench.
License. Our package (codes and datasets) is licensed under the MIT License. This license permits users to freely use, copy, modify, merge, publish, distribute, sublicense, and sell copies of the software, provided that the original copyright notice and permission notice are included in all copies or substantial portions of the software. The MIT License is widely accepted for its simplicity and permissive terms, ensuring ease of use and contribution to the codes and datasets. We bear all responsibility in case of violation of rights, etc, and confirmation of the data license.
Code Maintenance. We are committed to continuously updating our code and and actively addressing users’ issues and feedback. Additionally, we warmly welcome community contributions to enhance our library and benchmark algorithms. Nonetheless, we will enforce strict version control measures to ensure reproducibility throughout the maintenance process.
Appendix D Further Discussions
D.1 Related works.
Deep hypergraph learning (DHGL) have emerged as a powerful tool for mining the knowledge of higher-order interactions, which are pervasive in real-world complex systems and applications. Kim et al. (Kim et al., 2024b) recently present the first survey dedicated to HNNs, with an in-depth and step-by-step guide. The survey comprehensively reviews existing HNN architectures, training strategies, and applications, establishing a foundational understanding crucial for advancing the field of DHGL. However, systematic and comprehensive evaluations of HNN algorithms remain scarce. HyFER (Hwang et al., 2021) represents the first well-structured framework for fair HNN algorithm comparison, modularizing the training process into data, model, and task components. Moreover, to address the scalability problem that most existing implementations suffer from, HyFER is built on top of Deep Graph Library (DGL) (Wang et al., 2019), which is a highly-efficient open-sourced library for GNNs. THU-DH (Gao et al., 2022) is an open-source PyTorch-based toolbox designed for general HNNs. It supports various hypergraph pre-processing methods (e.g., sampling, expansion) and convolution operators, facilitating the evaluation of HNNs.
Despite the progress made by these benchmarks, their inherent limitations render it impossible to conduct a comprehensive and fair comparison across various HNN algorithms. Firstly, Existing benchmarks primarily focus on node- and edge-level datasets, while overlooking graph-level datasets. Moreover, the node-level datasets they employ are predominantly homophilic. For example, HyFER conducts benchmark evaluation using only four homophilic academic networks. Secondly, the algorithmic coverage of current benchmarks is limited. For instance, HyFER implements only three HNN models, while THU-DH supports only a subset of methods proposed before 2022, thus excluding many recent SOTA approaches. Thirdly, Existing benchmarks mainly focus on model effectiveness (e.g., accuracy), while overlooking other critical aspects such as efficiency and complexity. Additionally, robustness and fairness are essential for trustworthy HNN applications, yet are rarely evaluated. Finally, existing studies adopt inconsistent experimental protocols, including variations in hyperparameter settings, initialization procedures, and convergence criteria, which can affect reproducibility and the reliability of cross-study comparisons. To address these issues, we introduce DHG-Bench, a more comprehensive and fair benchmark for HNNs, aiming to advance the state of the art and provide deeper insights into the challenges and opportunities of deep hypergraph learning.
D.2 Limitations.
DHG-Bench has several limitations that we aim to address in our furture work.
Our current benchmark is limited to a specific set of tasks, which might not reflect the full potential and versatility of HNN models. By incorporate novel benchmark tasks (e.g., hyperedge disambiguation (Kim et al., 2023), edge-dependent node classification (Choe et al., 2023), and hyperlink prediction (Tang et al., 2025)), we can gain a deeper understanding of the progress in the field and provide insights into how different algorithms perform across diverse applications.
Due to resource constraints and the availability of implementations, we could not include some of the latest state-of-the-art HNNs in our benchmark. This might impact the comprehensiveness of our evaluation, as some promising methods are not represented. We aim to address this by continuously updating our package and incorporating these algorithms as they become available.
Our current benchmark primarily focuses on the semi-supervised learning paradigm. However, task-specific labels are often scarce in real-world scenarios, and deep models are prone to overfitting when trained with such limited supervision signals (Li et al., 2024a). In future evaluations, we plan to incorporate recently proposed hypergraph self-supervised learning methods (Wei et al., 2022; Lee & Shin, 2023; Li et al., 2024a; Kim et al., 2024a; Lee et al., 2024) to provide insights into HNNs’ robustness in label-scarce settings and their ability to learn transferable representations without heavy reliance on supervision.